OpenStudy (anonymous):

An unbiased coin is tossed four times. Find the probability of the given event. (Round your answer to four decimal places.) The coin lands heads more than once.

OpenStudy (dumbcow):

\[P(H>1) = 1 - P(H \le 1)\] so you need to find P(H=0) and P(H=1) use binomial thm \[P(H=0) = (.5)^{4}\] \[P(H=1) = 4*(.5)^{4}\]